Circular No. 5685

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1992 QB_{1}

Ephemeris continuation from the orbital elements on IAUC 5684:

1993TT α_{2000} δ_{2000} Δ r ϵ β V
Jan. 3 23^{h}56.^{m}04 − 0^{o}23.^{′}6 41.045 40.822 76.^{o}2 1.^{o}3 23.8
13 23 56.47 − 0 20.7 41.207 40.820 66.1 1.3 23.8
23 23 57.04 − 0 16.9 41.358 40.818 56.1 1.1 23.8

SATURN VI (TITAN) B. Bezard and A. Marten, Observatoire de Paris-Meudon; and G.

Paubert, Institut de Radio Astronomie Millimétrique, communicate: “We report the first detection of CH_{3}CN on Titan, using the IRAM 30-m radio telescope at Pico Veleta. The J=12-11 multiplet of CH_{3}CN around 220.7 GHz was observed at 0.1-MHz resolution on Dec. 25-27. During this run we also reobserved at the same spectral resolution the (24-23) line of HC_{3}N that we first detected in May. For both molecules the line width (FWHM) is ∼ 400 KHz, implying that the lines are formed in the doppler regime and originate from the upper stratosphere or mesosphere of Titan. A first analysis of these observations, based on radiative-transfer calculations, indicates that the disk-averaged column abundance of CH_{3}CN above 300 km is larger than that of HC_{3}N. These two millimeter observations impose stringent constraints on photochemical models of Titan.”

GRO J0422+32 B. A. Harmon and G. J. Fishman, NASA Marshall Space Flight Cen-

ter; and W. S. Paciesas, University of Alabama, Huntsville, report for the BATSE/CGRO Experiment Team: “Hard x-ray monitoring (cf. IAUC 5580, 5584, 5587, 5592, etc.) using earth occultation indicates that flux from GRO J0422+32 in the 20-300-keV energy band exhibits a secondary maximum similar to those of other x-ray novae. An upward turn in the hard x-ray flux began about Dec. 1, with the flux level at ∼ 150 mCrab, which then increased to a peak of ∼ 500 mCrab on Dec. 20. During Dec. 20-27 it dropped to about 400 mCrab. Prior to Dec. 1, the flux had been decreasing exponentially (time constant = 40.4 days) since the primary outburst in August. Fits to an optically-thin thermal bremsstrahlung model before and during the recent intensity increase yield average temperatures of 136 ± 8 keV (Nov. 24-Dec. 1) and 122 ± 4 keV (Dec. 15-22), respectively.”

1992 December 30 Brian G. Marsden